OpenStudy (anonymous):

Given the function f(x) = 3x, find the value of f−1(27).

OpenStudy (usukidoll):

replace f(x) with y y = 3x. Then solve for x meaning I need to have my x alone y/3 =x that should be my inverse y/3 = f^-1(y) if y = 27 then 9 = f^-1(y) it's been a long time since I have done this.. but I think I got it...?

OpenStudy (usukidoll):

sorry I'm super tired.. haven't slept well in a while.. no more school for me until late August so I can go late night on the pc

OpenStudy (anonymous):

well the choices are f−1(27) = 1 f−1(27) = 3 f−1(27) = 9 f−1(27) = 81

OpenStudy (anonymous):

@usukidoll

OpenStudy (usukidoll):

try the 3rd one

OpenStudy (anonymous):

i will. i have about 20 more questions on this test. ill let you know if its correct